Hii
I have a quetion related to number.
Adding 5 to four times of a number amounts to 53.find the number?

Found 2 solutions by stanbon, Aparesh Mondal:
Answer by stanbon(75887)

Adding 5 to four times of a number amounts to 53.find the number?
--------
4x + 5 = 53
-------
4x = 48
x = 12
-----------
Cheers,
Stan H.
----------------

Answer by Aparesh Mondal(1)

Let the number be x
then according to the question,
4x+5=53
or, 4x=53-5=48
or, x=48/4=12
therefore the number is 12